# Analysis of cohort stepped wedge cluster-randomized trials with non-ignorable dropout via joint modeling

### v1

@@ 4.8.2 Results: simulation study B, general time on treatment model
Figure 7 depicts the estimated bias for the treatment effects $\delta_{0},\delta_{1},\delta_{2},\delta_{3}$ on the longitudinal outcome over repetitions and across all simulation scenarios and models. These results are tabulated in Section C of the supplementary material. Overall, the joint model outperformed the linear mixed model, with no significant (or clinically meaningful) bias across the board; biases or larger magnitude, in absolute value, were observed for the later treatment effects (as expected, where the impact of informative data thinning is more pronounced). Scenarios that were not informative ($\omega_{1}=\omega_{2}=\log(1.0)$) showed no bias for either model; moreover, the mixed model performed well in scenarios with $\nu=0.0$ as well, despite the association parameters.
Bias for the period effects and ICCs followed the same patterns described in Section 4.8.1, with (1) the joint model outperforming the linear mixed model across the board when dropout was informative and (2) unbiased estimates when dropout was not informative.
Overall, compared to the results summarized in Section 4.8.1, the benefits of the joint modeling approach are more evident under the general time on treatment model. More results for the simulation under the general time on treatment model are included in Section C of the supplementary material available online.
Figure 7: Estimated bias for the treatment effects $\delta_{0},\delta_{1},\delta_{2},\delta_{3}$ on the longitudinal outcome with 95% confidence intervals based on Monte Carlo standard errors, general time on treatment model. Simulation scenarios based on the joint model.
@@ 5 Reanalysis of the ‘Frail Older Adults: Care in Transition’ trial
The ‘Frail Older Adults: Care in Transition’ (ACT) trial is a 24-month stepped wedge cluster randomized controlled trial that was conducted between May 2010 and March 2013 in 35 primary care practices in the Netherlands 45. The study included 1147 frail older adults, and aimed to study the impact of the Geriatric Care Model (GCM) on their quality of life; the study did not, however, show statistically significant beneficial effects of the GCM. One potential issue with the ACT study is dropout from the study: the study participants were frail older adults with an average age of 80 years at baseline, which are at high risk of dropping out of the study, e.g., because of death. In fact, after 24 months, more than 30% of study participants had dropped out.
@@
Thus, in this Section, we re-analyze data from the ACT trial using the joint modeling approach introduced in Section 3 to account for potentially informative dropout from the study; we also include results from the standard linear mixed model analysis for comparison purposes. Specifically, we focus on two longitudinal outcomes: the mental health component score (MCS) and physical health component score (PCS) of the 12-item Short Form questionnaire (SF-12), which quantify quality of life. We fit joint models assuming a constant intervention effect or a general time on treatment effect, equivalent to those introduced in Equations (3) and (5), respectively, and adjust the longitudinal submodels for age, sex, and baseline variables on which the allocation groups differed at baseline (educational level, region and frailty index score), as in Hoogendijk et al. 45 The survival submodel included time-varying treatment as a covariate and assumed the shared random effects association structure. Comparable linear mixed models were fit by ignoring the time-to-event component.
@@
The estimated treatment effect parameters are depicted in Figure 8: despite explicitly modeling the dropout process, we still estimated a small and not clinically meaningful association between the GCM and quality of life as measured by SF-12. Moreover, differences between the joint and linear mixed models were minor and did not change the interpretation of the results. Estimated period effects, variance components, and, therefore, ICC values were also close when comparing the joint and linear mixed models, as illustrated in Appendix D of the Supplementary Material available online.
@@
These results suggest that dropout from the study in ACT is likely not strongly informative; we can test this assumption by studying the estimated coefficients of the survival submodel, which are listed in Table 1. Specifically, the estimated parameter $\nu$ is close to zero for both models, showing that there is no direct effect of treatment on the hazard of dropout. Moreover, the estimated $\omega_{1}$ parameters are also close to zero and with large estimated standard errors, thus suggesting no association between the cluster-specific random intercept and dropout; conversely, the estimated $\omega_{2}$ parameters suggest a significant association between the subject-specific random intercept and dropout but of small magnitude (relative to the size of the estimated variances, 37.46 and 37.51 for the constant treatment effect and general time on treatment joint model, respectively; Appendix D, Supplementary Material) for the SF-12 MCS outcome, and a not clinically meaningful association (once again) for the SF-12 PCS outcome. Overall, the estimated values of $\omega_{1},\omega_{2}$ suggest that either there is no differential dropout, e.g., between cluster or study subjects, or that this effect is of small magnitude and with negligible clinical relevance.
@@
These results provide a useful sensitivity analysis for the ACT trial regarding the issue of possibly informative dropout. Our analyses suggest that the estimated treatment effects from the mixed model are robust to violations of this assumption (in these settings).
@@
   $\nu$0.43 (-0.71, -0.15) -0.43 (-0.71, -0.14)
   $\omega_{1}$0.48 (-2.55, 1.58) -1.14 (-12.47, 10.18)
   $\omega_{2}$0.02 (-0.05, -0.00) -0.02 (-0.05, -0.00)
   $\log(\lambda)$1.47 (-1.61, -1.33) -1.47 (-1.61, -1.33)
   $\log(p)$0.06 (-0.17, 0.05) -0.06 (-0.17, 0.05)
@@
   $\nu$0.36 (-0.62, -0.10) -0.36 (-0.61, -0.10)
   $\omega_{1}$0.11 (-0.32, 0.09) -0.11 (-0.32, 0.10)
   $\omega_{2}$0.01 (-0.04, 0.01) -0.01 (-0.04, 0.01)
   $\log(\lambda)$1.47 (-1.60, -1.34) -1.47 (-1.60, -1.34)
   $\log(p)$0.08 (-0.19, 0.03) -0.08 (-0.19, 0.03)
@@ 6 Extensions of the joint model to other outcome types and association structures
The joint modeling framework can be extended in a variety of ways. In this Section, we introduce and discuss two of these: the extension to generalized outcome types (such as binary or count) and the extension to additional, more interpretable association structures between the longitudinal and time-to-event submodels.
@@
Similar extensions can be defined for the general time on treatment models of Equations (2) and (5); these are omitted from this Section for simplicity. Note that these generalized joint models can be, once again, fitted using the gsem command in Stata 34; every built-in distribution for the longitudinal outcome is supported within this framework, with a certain amount of additional computational complexity depending on the distribution of choice.
For the second proposed extension, recall that in every joint model introduced so far, we assumed the shared random effects parametrization for the association structure between the longitudinal and dropout submodels. The interpretation of the estimated association parameters reflects the change in the log hazard of dropout for a unit change in the deviation of each cluster/subject from the population mean; note that this association structure is time-independent, as the latent random intercept values do not change over time. However, this parameterization may be challenging to interpret for more elaborate structures; thus, alternative association structures have been proposed in the literature 46, 47.
One such alternative association structure is the so-called expected value association structure, where we link the two submodels via the expected value of the longitudinal outcome at a certain point in time $t$. For instance, the joint model of Equation (3) could be re-written as:
@@
where $E[Y_{ijk}(t)]$ denotes the expected value of the outcome $Y$ at time $t$. The association parameter $\omega$ can now be interpreted as the change in log hazard of dropout for a unit change in the expected value of the outcome $Y$ at time $t$. Note that the expected outcome value changes over time, which introduces additional computational complexity in the estimation process. Alternative association structures are discussed elsewhere and include, among others, the rate of change in the longitudinal outcome at time $t$ or the (possibly weighted) cumulative effect 47. Alternative association structures are not yet available within Stata’s gsem command and would require ad-hoc software development to accommodate all the additional complexities of stepped wedge trials (such as the hierarchical structure and the time-varying exogenous treatment assignment) before being usable in practice.
@@ 7 Discussion
We introduced a joint longitudinal-survival model that can be used in the settings of closed-cohort stepped wedge to account for non-ignorable dropout. The proposed model can be used with both constant and general time on treatment effect parametrizations, can accommodate time-varying exogenous covariates in the dropout submodel (such as treatment), and can accommodate different ways of linking the submodels (e.g., the user-defined form of $\omega$ in Equation (3)). Moreover, we provide annotated technical guidance and statistical code for readers to apply the methodology in practice, including synthetic datasets — this is openly available on GitHub at https://github.com/ellessenne/swjm.
@@
Using Monte Carlo simulation and realistic data-generating mechanisms, we showed that the joint model outperformed the standard linear mixed effects model in a variety of scenarios with informative dropout. Specifically, the joint model yielded unbiased treatment and period effect estimates, with the mixed model performing significantly worse in estimating period effects and treatment effects under the general time on treatment parametrization. Moreover, the mixed model yielded significant bias in the estimated variance components and, therefore, the ICCs. The mixed model was, however, relatively robust in some of the scenarios with informative dropout, with unbiased treatment effect estimates at the cost of bias in the estimated variance components, which were inflated to accommodate the additional heterogeneity due to informative dropout. Interestingly, fitting the joint model in scenarios where dropout was not informative did not introduce any bias in the analysis. This highlights that fitting the joint model is a viable strategy to confirm the results of a mixed model when it is unclear whether dropout from the study is informative or not, as a possible sensitivity analysis.
@@
This manuscript has some strengths and limitations. Among the strengths, our Monte Carlo simulation study is structured to follow the ADEMP framework 37 and used realistic data-generating mechanisms informed by the ACT study, which was re-analyzed in Section 5. The joint modeling approach can be implemented using off-the-shelf statistical software and can be easily extended to accommodate other variants of the shared random effects association structure, different treatment effects parametrizations (e.g., the linear time-on-treatment and delayed treatment effect parametrizations discussed by Li et al. 3), and baseline hazard functions for the dropout submodel. These extensions can be easily implemented using gsem in Stata. We also developed an algorithm for simulating data from the joint model in the settings of stepped wedge CRTs, which is not trivial; we implemented this in an R package that is available on GitHub at https://github.com/RedDoorAnalytics/simswjm, with statistical code to replicate our simulation study and implement the joint model in practice also available on GitHub at https://github.com/ellessenne/swjm. Among the limitations of our study, the data-generating mechanisms are based on the joint model, therefore the good performance of the joint model is not entirely surprising 48. Nonetheless, the results of this manuscript are in agreement with the current literature on joint longitudinal-survival modeling 17, 18. Thomadakis et al. showed that comparable joint longitudinal-survival models could be biased in certain MAR scenarios, but we did not observe such behavior — likely due to the non-informative scenarios reducing to MCAR, where joint models are expected to estimate unbiased results 49, 50. The results of the simulation study can only be generalized to settings close to the assumed data-generating mechanisms. Implementing and reporting a larger number of scenarios would likely become unwieldy, and further ad-hoc studies are therefore warranted. Next, we modeled the dropout process in continuous time (compared to, e.g., discrete-time models as in Wang and Chinchilli 23), therefore we need to assume that dropout times are known nearly exactly. However, extensions to the methodology to accommodate interval censoring are possible. Of note, our methods can be directly applied to the individually-randomized stepped wedge design 51, which is a recent variation of the traditional individually-randomized parallel-arm design by including staggered treatment assignment for efficiency improvement. In that case, the appropriate joint model is a special case of the model of Equation (3) after removing the cluster-level random effect $\alpha_{i}$, and we provide sample code to implement this procedure in the GitHub repository. Finally, some of the extensions discussed throughout the manuscript and in Section 6 require further software developments before they can be used in practice. This is also left for future work.
@@
This work can also be further extended beyond what was introduced and discussed in Section 6. For instance, we currently use a frequentist approach with maximum likelihood estimation: a natural extension is to frame the joint model within a Bayesian framework, to allow the incorporation of prior information (e.g., from previous trials). Moreover, we do not consider that some of the outcome measures are non-directly measurable, i.e., with studies using a set of observed indicators from questionnaires or measurement scales to evaluate an unobservable latent variable. Saulnier et al. propose a latent process joint model to accommodate such outcomes, which could be combined with the methodology of this manuscript for the settings of stepped wedge trials 52. Finally, there exist more complicated random-effects structures for modeling longitudinal outcomes in closed-cohort stepped wedge designs, such as those that incorporate a random cluster-by-period effect and further allow for discrete-time correlation decay 53, 3. These extensions will inevitably introduce additional computational complexity under the joint modeling framework and will be important directions for future research.
@@
In conclusion, we described joint modeling methodology that can be used in the settings of stepped wedge cluster randomized trials when dropout from the study is suspected to be informative, and we showed that the joint modeling approach reduced bias across the board compared to standard linear mixed effects models that ignore the dropout process. Annotated statistical code is provided for readers to apply this methodology in practice.
@@
A. G.: conceptualization, methodology, software, writing - original draft, visualization, formal analysis, data curation, project administration. F. L.: conceptualization, methodology, writing - review & editing, supervision, funding acquisition. E. O. H: resources (provision of ACT data), writing - review & editing. M. J. C.: methodology, software, writing - review & editing, supervision. M. O. H.: conceptualization, methodology, writing - review & editing, supervision, funding acquisition.
@@
Additional results for the simulation studies reported in Section 4.8 and for the applied example reported in Section 5 are included in the Supplementary Material, available online on the journal website. Statistical code to replicate the simulation study and an R package implementing the simulation algorithm are available on GitHub at https://github.com/ellessenne/swjm and https://github.com/RedDoorAnalytics/simswjm, respectively.

### v3

@@ 5.8.2 Results: simulation study B, general time on treatment model
Figure 7 depicts the estimated bias for the treatment effects $\delta_{0},\delta_{1},\delta_{2},\delta_{3}$ on the longitudinal outcome over repetitions and across all simulation scenarios and models. These results are tabulated in Appendix B of the supplementary material. Overall, the joint model outperformed the linear mixed model, with no apparent bias across the board; biases in larger magnitude were observed for the treatment effects under higher level of exposure time; this is as expected because the effective sample size for estimating those parameters is smaller and the impact of informative data thinning due to dropout is more pronounced. Scenarios that were not informative ($\omega_{1}=\omega_{2}=\log(1.0)$) showed no bias for either model; moreover, the mixed model performed well in scenarios with $\nu=0.0$ as well, across values of the association parameters. Coverage probability was close to nominal for both models.
Bias for the period effects and ICCs followed the same patterns described in Section 5.8.1. That is, (1) the joint model outperformed the linear mixed model across the board when dropout was informative and (2) both models returned unbiased estimates when dropout was not informative.
Overall, compared to the results summarized in Section 5.8.1, the benefits of the joint modeling approach are more evident under the general time on treatment model. More results for the simulation under the general time on treatment model are included in Appendix B of the supplementary material available online.
Figure 7: Estimated bias for the treatment effects $\delta_{0},\delta_{1},\delta_{2},\delta_{3}$ on the longitudinal outcome with 95% confidence intervals based on Monte Carlo standard errors, general time on treatment model.
5.8.3 Results: additional simulation scenarios for both model formulations
Estimated bias for the treatment effect parameter(s) and the additional (non neutral) simulation scenarios are depicted in Figure 8. The linear mixed model yielded biased estimates under every additional scenario, with greater bias in the scenario with fewer clusters per intervention sequence. Superiority of the joint model was confirmed under every one of these additional scenarios with informative dropout: additional results are included in Appendix B of the supplementary material available online.
Figure 8: Estimated bias for the treatment effect $\delta$ on the longitudinal outcome with 95% confidence intervals based on Monte Carlo standard errors, additional simulation scenarios for the constant intervention and general time on treatment models. Note that $i$ denotes the total number of clusters, and $N$ denotes the number of participants per cluster.
Finally, both models yielded biased estimates in the neutral simulation scenario: this is likely because both models are misspecified by the simulation design in such setting. Nonetheless, the joint model still outperformed the linear mixed model across the board, with less bias for all model parameters (including ICCs and variance components). These results are summarized in Appendix B of the supplementary material.
@@ 6 Reanalysis of the ‘Frail Older Adults: Care in Transition’ trial
The ‘Frail Older Adults: Care in Transition’ (ACT) trial is a 24-month stepped wedge cluster randomized controlled trial that was conducted between May 2010 and March 2013 in 35 primary care practices in the Netherlands 54. The study included 1147 frail older adults, and aimed to study the impact of the Geriatric Care Model (GCM) on their quality of life; the study did not, however, show statistically significant beneficial effects of the GCM. One potential issue with the ACT study is dropout from the study: the study participants were frail older adults with an average age of 80 years at baseline, which are at high risk of dropping out of the study, e.g., because of death. In fact, after 24 months, more than 30% of study participants had dropped out.
@@
Thus, in this Section, we re-analyze data from the ACT trial using the joint modeling approach introduced in Section 4 to account for potentially informative dropout from the study; we also include results from the standard linear mixed model analysis for comparison purposes. Specifically, we focus on two longitudinal outcomes: the mental health component score (MCS) and physical health component score (PCS) of the 12-item Short Form questionnaire (SF-12), which quantify quality of life. We fit joint models assuming a constant intervention effect or a general time on treatment effect, equivalent to those introduced in Equations (3) and (5), respectively, and adjust the longitudinal submodels for age, sex, and baseline variables on which the allocation groups differed at baseline (educational level, region and frailty index score), as in Hoogendijk et al. 54 The survival submodel included time-varying treatment as a covariate and assumed the shared random effects association structure. Comparable linear mixed models were fit by ignoring the time-to-event component.
@@
The estimated treatment effect parameters are depicted in Figure 9: despite explicitly modeling the dropout process, we still estimated a small but not clinically meaningful association between the GCM and quality of life as measured by SF-12. This estimated association was not statistically significant. Overall, we observed treatment effect estimates closer to the null under the linear mixed model as compared to the joint model, but differences between the two were minor and did not change the interpretation of the results. Estimated period effects, variance components, and, therefore, ICC values were also close when comparing the joint and linear mixed models, as illustrated in Appendix D of the Supplementary Material available online.
@@
Due to the minor differences between estimates from the two models, we infer that dropout from the study in ACT is likely not strongly informative; we can test this assumption by studying the estimated coefficients of the survival submodel, which are listed in Table 1. Specifically, the estimated parameter $\nu$ suggests a lower dropout rate for treated individuals, highlighting a direct effect of treatment on the hazard of dropout. Moreover, the estimated $\omega_{1}$ parameters are close to zero and with large estimated standard errors, thus likely suggesting no association between the cluster-specific random intercept and dropout. In contrast, the estimated $\omega_{2}$ parameter suggests a significant association between the subject-specific random intercept and dropout but of small magnitude (relative to the size of the estimated variances; Appendix D, Supplementary Material) for the SF-12 MCS outcome, where the estimated $\omega_{2}$ parameter is approximately zero for the SF-12 PCS outcome. Overall, the estimated values of $\omega_{1},\omega_{2}$ suggest that either there is no differential dropout, e.g., between cluster or study subjects, or that this effect is of small magnitude leading to only minor impact on the final treatment effect estimates on the SF-12 outcomes.
@@
In this example, the use of the joint modeling provides a useful sensitivity analysis for the ACT trial regarding the issue of possibly informative dropout. In conclusion, our analyses suggest that the estimated treatment effects from the mixed model are only slight weaker but generally close to those from the joint model, likely because the dropout is not strongly informative.
@@
   $\nu$0.415 (-0.682, -0.149) -0.416 (-0.682, -0.149)
   $\omega_{1}$0.305 (-0.831, 0.221) -0.303 (-0.822, 0.216)
   $\omega_{2}$0.035 (-0.067, -0.002) -0.035 (-0.067, -0.002)
   $\log(\lambda)$1.475 (-1.607, -1.344) -1.475 (-1.607, -1.343)
   $\log(p)$0.065 (-0.174, 0.044) -0.064 (-0.173, 0.044)
@@
   $\nu$0.382 (-0.641, -0.123) -0.387 (-0.647, -0.126)
   $\omega_{1}$ 0.061 (-0.074, 0.197) 0.066 (-0.067, 0.200)
   $\omega_{2}$0.000 (-0.040, 0.040) -0.001 (-0.041, 0.040)
   $\log(\lambda)$1.452 (-1.580, -1.324) -1.451 (-1.579, -1.323)
   $\log(p)$0.076 (-0.184, 0.032) -0.075 (-0.183, 0.033)
@@ 7 Extensions of the joint model to other outcome types and association structures
The joint modeling framework can be extended in a variety of ways. In this Section, we briefly introduce and discuss two of these: the extension to generalized outcome types (such as binary or count) and the extension to additional, more interpretable association structures between the longitudinal and time-to-event submodels.
@@
Similar extensions can be defined for the general time on treatment models of Equations (2) and (5); these are omitted from this Section for simplicity. Note that these generalized joint models can be fitted using the gsem command in Stata; 42 every built-in distribution for the longitudinal outcome is supported within this framework, with a certain amount of additional computational complexity depending on the distribution of choice. Example statistical code to fit these joint models is included on GitHub at https://github.com/RedDoorAnalytics/simswjm.
For the second proposed extension, recall that in every joint model introduced so far, we assumed the shared random effects parametrization for the association structure between the longitudinal and dropout submodels. The interpretation of the estimated association parameters reflects the change in the log hazard of dropout for a unit change in the deviation of each cluster/subject from the population mean; note that this association structure is time-independent, as the latent random intercept values do not change over time. Alternative, time-dependent association structures have been proposed in the literature. 30, 55 One such alternative association structure is the so-called expected value association structure, where we link the two submodels via the expected value of the longitudinal outcome at a certain point in time $t$. For instance, the joint model of Equation (3) could be re-written as:
@@
where $E[Y_{ijk}(t)]$ denotes the expected value of the outcome $Y$ at time $t$. The association parameter $\omega$ can now be interpreted as the change in log hazard of dropout for a unit change in the expected value of the outcome $Y$ at time $t$. Note that the expected outcome value changes over time, which introduces additional computational complexity in the estimation process. Alternative association structures are discussed elsewhere and include, among others, the rate of change in the longitudinal outcome at time $t$ or the (possibly weighted) cumulative effect. 55 Alternative association structures are not yet available within Stata’s gsem command and would require ad-hoc software development to accommodate all the additional complexities of stepped wedge trials (such as the hierarchical structure and the time-varying exogenous treatment assignment) before being usable in practice.
@@ 8 Discussion
We introduced a joint longitudinal-survival model that can be used in the settings of closed-cohort stepped wedge CRTs to account for non-ignorable dropout. The proposed model can be used with both constant and general time on treatment effect parametrizations, can accommodate time-varying exogenous covariates in the dropout submodel (such as treatment), and can accommodate different ways of linking the submodels (e.g., the user-defined form of $\omega$ in Equation (3)). Moreover, we provide annotated technical guidance and statistical code for readers to apply the methodology in practice, including synthetic datasets — this is openly available on GitHub at https://github.com/ellessenne/swjm and archived on Zenodo as well. 52 Usual model-building strategies apply, e.g., in terms of functional form of covariates and interaction terms, baseline hazard distributions, random effects specification. Fit of the joint model with the observed data can be assessed by comparing different model formulations using, e.g., information criteria (such as the Akaike and Bayesian Information Criteria 56, 57, 58) or by comparing predicted and observed values. Robustness of the random effects assumptions increases with the number of observations per cluster; 30 in small sample size settings, $t$-distributed random effects could be used instead, as long as the software implementation supports it. 59
Moreover, we developed an algorithm to simulate data from stepped wedge CRTs with or without informative dropout, which can be useful beyond the settings of this manuscript, e.g., to estimate the required sample size for a future study using a simulation-based approach. The code is openly available for other to re-use and apply at https://github.com/RedDoorAnalytics/simswjm.
@@
Using Monte Carlo simulation and realistic data-generating mechanisms, we showed that the joint model outperformed the standard linear mixed-effects model in a variety of scenarios with informative dropout. Specifically, the joint model yielded unbiased treatment and period effect estimates, with the mixed model performing significantly worse in estimating period effects and treatment effects under the general time on treatment parametrization. Moreover, the mixed model yielded significant bias in the estimated variance components and, therefore, the ICCs. The mixed model was, however, relatively robust in some of the scenarios with informative dropout, with unbiased treatment effect estimates at the cost of bias in the estimated variance components, which were inflated to accommodate the additional heterogeneity due to informative dropout. Interestingly, fitting the joint model in scenarios where dropout was not informative did not introduce any bias in the analysis. This highlights that fitting the joint model is a viable strategy to confirm the results of a mixed model when it is unclear whether dropout from the study is informative or not, as a possible sensitivity analysis.
@@
This manuscript has some strengths and limitations. Among the strengths, our Monte Carlo simulation study is structured to follow the ADEMP framework and used realistic data-generating mechanisms informed by the ACT study, which was re-analyzed in Section 6. The joint modeling approach can be implemented using off-the-shelf statistical software and can be easily extended to accommodate other variants of the shared random effects association structure, different treatment effects parametrizations (e.g., the linear time-on-treatment and delayed treatment effect parametrizations discussed by Li et al. 3), and baseline hazard functions for the dropout submodel. These extensions can be easily implemented using gsem in Stata. Our approach can accommodate time-varying, exogenous covariates in the survival submodel, which is not trivial; as we demonstrate how to fit this with off-the-shelf statistical software, this can be easily extended to other settings, study designs, and disease areas. We also developed an algorithm for simulating data from the joint model in the settings of stepped wedge CRTs; we implemented this in an R package that is available on GitHub at https://github.com/RedDoorAnalytics/simswjm, with statistical code to replicate our simulation study and implement the joint model in practice also available on GitHub at https://github.com/ellessenne/swjm.
Among the limitations of our study, the main data-generating mechanisms are based on the joint model, therefore the good performance of the joint model is somewhat expected. 60 Nonetheless, superiority of the joint model was confirmed in a neutral simulation scenario that we tested, and the results of this manuscript are in agreement with the current literature on joint longitudinal-survival modeling, while providing new evidence in the settings of hierarchical models with more that one level of nesting. 18, 19 Thomadakis et al. showed that comparable joint longitudinal-survival models could be biased in certain MAR scenarios, but we did not observe such behavior — likely due to the non-informative scenarios reducing to MCAR, where joint models are expected to estimate unbiased results. 61, 62 The results of the simulation study can only be generalized to settings close to the assumed data-generating mechanisms. Implementing and reporting a larger number of scenarios would likely become unwieldy, and further simulation studies are therefore warranted. Next, we modeled the dropout process in continuous time (compared to, e.g., discrete-time models as in Wang and Chinchilli 24), therefore we need to assume that dropout times are known nearly exactly. Of note, our methods can be directly applied to the individually-randomized stepped wedge design, 63 which is a recent variation of the traditional individually-randomized parallel-arm design by including staggered treatment assignment for efficiency improvement. In that case, the appropriate joint model is a special case of the model of Equation (3) after removing the cluster-level random effect $\alpha_{i}$, and we provide sample code to implement this procedure in the GitHub repository. Finally, some of the extensions discussed throughout the manuscript and in Section 7 require further software developments before they can be used in practice. This is also left for future work.
@@
This work can also be further extended beyond what was introduced and discussed in Section 7. For instance, we currently use a frequentist approach with maximum likelihood estimation: a natural extension is to frame the joint model within a Bayesian framework, to allow the incorporation of prior information (e.g., from previous trials). Moreover, we do not consider that some of the outcome measures are non-directly measurable, i.e., with studies using a set of observed indicators from questionnaires or measurement scales to evaluate an unobservable latent variable. Saulnier et al. propose a latent process joint model to accommodate such outcomes, which could be combined with the methodology of this manuscript for the settings of stepped wedge trials. 64 Finally, there exist more complicated random-effects structures for modeling longitudinal outcomes in closed-cohort stepped wedge designs, such as those that incorporate a random cluster-by-period effect and further allow for discrete-time correlation decay. 65, 3 These extensions will inevitably introduce additional computational complexity under the joint modeling framework and will be important directions for future research.
@@
In conclusion, we described joint modeling methodology that can be used in the settings of stepped wedge cluster randomized trials when dropout from the study is suspected to be informative, and we showed that the joint modeling approach reduced bias across the board compared to standard linear mixed-effects models that ignore the dropout process. Annotated statistical code is provided for readers to apply this methodology in practice.
@@
A. G.: conceptualization, methodology, software, writing - original draft, writing - review & editing, visualization, formal analysis, data curation, project administration. F. L.: conceptualization, methodology, writing - review & editing, supervision, funding acquisition. E. O. H: resources (provision of ACT data), writing - review & editing. M. J. C.: methodology, software, writing - review & editing, supervision. M. O. H.: conceptualization, methodology, writing - review & editing, supervision, funding acquisition.
@@
Additional results for the simulation studies reported in Section 5.8 and for the applied example reported in Section 6 are included in the Supplementary Material, available online on the journal website. Statistical code to replicate the simulation study and an R package implementing the simulation algorithms are available on GitHub at https://github.com/ellessenne/swjm and https://github.com/RedDoorAnalytics/simswjm, respectively. Replication and supporting code is archived on Zenodo as well. 52